40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9593     Entry time: Mon Feb 3 23:31:33 2014
Author: Manasa 
Type: Update 
Category: General 
Subject: Alignment update / Y arm locked 

[EricQ, Manasa, Koji]

We measured the spot positions on the MC mirrors and redid the MC alignment by only touching the MC mirror sliders. Now all the MC spots are <1mm away from the center.

We opened the ITMY and ETMY chambers to align the green to the arm. The green was already centered on the ITMY. We went back and forth to recenter the green on the ETMY and ITMY (This was done by moving the test masses in pitch and yaw only without touching the green pointing) until we saw green flashes in higher order modes. At this point we found the IR was also centered on the ETMY and a little low in pitch on ITMY. But we could see IR flashes on the ITMYF camera. We put back the light doors and did the rest of the alignment using the pitch and yaw sliders.

When the flashes were as high as 0.05, we started seeing small lock stretches. Playing around with the gain and tweaking the alignment, we could lock the Y arm in TEM00 for IR and also run the ASS. The green also locked to the arm in 00 mode at this point. We aligned the BS to get a good AS view on the camera. ITMX was tweaked to get good michelson.

ELOG V3.1.3-